# DisasterM3: A Remote Sensing Vision-Language Dataset for Disaster Damage Assessment and Response

### v1

@@ 4.1 Comparative Results
Domain gap for disaster scenarios. Tab. 2 presents performance evaluations on optical-optical settings for QA tasks. As a traditional VLM, LLaVA-1.5 exhibited significant limitations when processing disaster scenes due to the domain gap. By leveraging extensive multi-modal pretraining datasets and implementing the AnyRes architecture, LLaVA-OV demonstrates enhancements in both accuracy and multi-image processing capabilities. As efficient Mixture-of-Experts (MoE) VLMs, Kimmi-VL-Think exceeds Kimmi-VL-Instruct in mathematical counting tasks (BDC, DRE). However, the non-negligible domain gap limits their application on complex tasks, particularly degrading performance to near-random levels on the ORR task. This motivated our development of the DisasterM3 dataset, which identifies performance gaps through the Bench set while providing complementary training data via the Instruct set.
@@
LLaVA-1.5-7B [19] 12.1 4.2 - - - - 20.0 - - - - - - - -
LLaVA-OV-7B [17] 24.5 16.3 53.5 3.7 26.4 24.2 22.7 1.66 1.50 1.53 1.93 2.30 3.01 2.08 1.81
Kimi-VL-A3B-Instruct [35] 25.6 28.9 66.3 4.0 20.4 15.0 18.9 1.69 1.53 1.72 1.81 2.67 3.57 2.40 2.05
Kimi-VL-A3B-Think [35] 26.7 27.0 51.6 7.4 24.4 25.4 24.4 1.61 1.39 1.68 1.75 2.61 3.35 2.34 2.15
InternVL3-8B [53] 31.3 39.6 53.5 4.0 30.3 24.1 36.2 1.96 1.88 1.92 2.09 2.75 3.52 2.53 2.21
InternVL3-14B [53] 35.7 42.5 62.0 4.9 27.4 23.6 54.1 2.08 2.01 2.01 2.22 2.86 3.67 2.62 2.29
InternVL3-78B [53] 39.3 43.5 72.5 5.3 29.4 28.7 56.1 2.79 2.74 2.75 2.89 2.90 3.64 2.64 2.43
Qwen2.5-VL-3B [3] 26.2 30.8 56.1 5.7 29.9 21.2 13.8 1.00 0.83 1.05 1.12 2.15 2.98 1.77 1.71
Qwen2.5-VL-7B [3] 31.2 28.3 66.6 4.7 34.2 29.3 23.9 1.75 1.69 1.71 1.85 1.95 2.53 1.83 1.49
Qwen2.5-VL-32B [3] 35.3 36.7 54.7 11.6 33.2 30.9 44.8 1.55 1.42 1.52 1.72 2.96 3.63 2.71 2.55
Qwen2.5-VL-72B [3] 40.5 47.0 74.8 6.8 34.8 28.9 50.8 2.01 1.99 2.00 2.05 2.92 3.79 2.70 2.27
GeoChat-7B [14] 10.7 6.1 - - - - 15.3 - - - - - - - -
TeoChat-7B [13] 23.0 6.9 64.9 2.0 22.5 23.3 18.2 1.77 1.61 1.74 1.96 1.95 2.59 1.77 1.49
EarthDial-4B [34] 22.9 10.6 58.1 3.2 30.2 20.8 14.5 1.53 1.22 1.64 1.73 2.42 3.21 2.08 1.98
@@
Qwen2.5-VL-7B [3] 40.4 37.7 83.6 21.5 34.3 29.4 36.2 3.90 3.76 3.53 4.41 3.11 3.73 2.88 2.73
$\Delta$uparrow9.2 $\uparrow$9.4 $\uparrow$17.0 $\uparrow$16.8 $\uparrow$0.1 $\uparrow$0.1 $\uparrow$12.3 $\uparrow$2.15 $\uparrow$2.07 $\uparrow$1.82 $\uparrow$2.56 $\uparrow$1.26 $\uparrow$1.20 $\uparrow$1.83 $\uparrow$1.24
InternVL3-8B [53] 41.7 42.6 79.3 23.9 29.1 24.9 50.6 3.83 3.69 3.49 4.32 3.31 3.92 3.10 2.90
$\Delta$uparrow10.4 $\uparrow$3.0 $\uparrow$25.8 $\uparrow$19.9 $\downarrow$-1.2 $\uparrow$0.8 $\uparrow$14.4 $\uparrow$1.87 $\uparrow$1.81 $\uparrow$1.57 $\uparrow$2.23 $\uparrow$0.56 $\uparrow$0.40 $\uparrow$0.57 $\uparrow$0.69
Larger VLMs achieve higher performances. By scaling up LLMs, InternVL3 and Qwen2.5-VL series demonstrate consistent trends that larger LLMs achieve superior performances, confirming established scaling laws observed in general-domain applications. The commercial models, i.e., GPT4o and GPT4.1, showcase competitive performances across all tasks due to their massive corpus.
@@
Fine-tuned models improve comprehensively. By fine-tuning on DisasterM3 Instruct set, the performances of Qwen2.5-VL and InternVL3 have been significantly improved, narrowing the domain gap. Disaster-specific terminology integration during training significantly enhances report generation quality, resulting in more reasonable and professional reports. However, for building damage counting (BDC) task, the fine-tuned InternVL3 exhibits unexpected performance degradation due to overfitting, and we perform detailed analysis in §4. In the future, object sensitive module [52] and numerical enhanced optimization [41] could be explored for model development.
@@
Underrepresentation for SAR images. Disasters are often accompanied by extreme weather, with clouds and rain blocking optical sensors. In this case, the active imaging method SAR can penetrate clouds and fog to obtain accurate surface information. Fig. 7 shows the VLMs’ performances evaluated on paired optical-SAR images. Due to the reduced semantics compared to optical imagery and underrepresentation in generic VLM, the performance using post-SAR images yielded substantially diminished performance across all evaluation tasks. In this scenario, as demonstrated by the Qwen2.5-VL series, increasing LLM size fails to yield stable improvements. The finetuned models alleviate the multi-sensor feature alignment, significantly improving the performances on all tasks. Since there still exists a huge gap compared to the optical-optical setting, more multi-modal pretraining [10] and alignment strategies could be investigated [40] for further improvement.
Table 3: Benchmarking results of referring segmentation VLMs on DisasterM3 Bench set. The accuracies of damage-levels (buildings and roads) are merged for simplification.
@@
PSALM-1.3B [50] 9.7 6.3 2.6 10.2 16.3 8.1 8.8 5.1 11.1
HyperSeg-3B [44] 16.6 14.5 7.5 17.0 25.4 8.8 10.3 4.5 13.1
LISA-7B [15] 27.5 22.1 11.9 25.0 45.6 10.9 12.7 6.0 15.7
GeoPixel-7B [32] 14.3 14.2 8.5 18.1 16.2 4.0 5.1 1.8 6.2
@@
LISA-7B [15] 44.8 43.7 27.6 41.2 65.5 28.2 29.6 21.5 34.9
$\Delta$uparrow17.3 $\uparrow$21.6 $\uparrow$15.7 $\uparrow$16.2 $\uparrow$19.9 $\uparrow$17.3 $\uparrow$16.9 $\uparrow$15.5 $\uparrow$19.2
PSALM-1.3B [50] 50.5 44.5 30.5 49.1 71.9 31.8 35.2 24.3 39.3
$\Delta$uparrow40.8 $\uparrow$38.2 $\uparrow$27.9 $\uparrow$38.9 $\uparrow$55.6 $\uparrow$23.7 $\uparrow$26.4 $\uparrow$19.2 $\uparrow$28.2
@@ 4.2 Detailed Analysis
LLaVA-OV [17] 21.2 23.6 17.1 19.2 22.8 25.4 18.1 19.5 23.2 23.5 19.9
Kimi-VL-A3B-Instruct [35] 19.9 22.2 17.5 20.1 16.3 26.3 13.2 19.2 21.9 23.5 18.4
Kimi-VL-A3B-Think [35] 22.0 26.4 19.6 21.0 17.4 26.9 16.9 22.6 21.8 25.0 22.3
InternVL3-8B [53] 27.5 41.7 22.2 24.4 21.7 33.0 20.9 28.0 27.3 28.8 27.0
InternVL3-14B [53] 30.0 48.6 22.7 26.6 27.2 33.7 21.1 27.3 29.9 31.3 31.5
InternVL3-78B [53] 31.8 48.6 26.3 27.9 25.0 37.1 25.6 32.0 31.7 32.9 30.9
Qwen2.5-VL-3B [3] 24.5 26.4 19.5 21.9 27.5 32.1 17.9 24.2 24.2 29.6 21.8
Qwen2.5-VL-7B [3] 25.6 34.3 21.0 24.9 17.3 33.7 16.7 28.3 27.8 25.6 25.9
Qwen2.5-VL-32B [3] 31.0 50.0 26.4 27.1 26.6 35.7 23.9 32.8 29.4 30.8 27.7
Qwen2.5-VL-72B [3] 31.8 47.2 25.0 31.1 19.0 39.0 24.0 33.4 34.0 33.9 31.2
GPT4o [12] 30.7 52.8 24.8 25.7 19.6 33.7 25.6 28.0 29.7 35.1 32.0
GPT4.1 [12] 32.4 51.4 26.9 26.7 21.7 35.2 27.7 28.5 33.4 35.8 36.5
@@
Qwen2.5-VL-7B [3] 32.9 41.7 26.5 30.7 27.7 40.3 22.3 33.0 34.0 41.8 31.1
$\Delta$uparrow7.3 $\uparrow$7.4 $\uparrow$5.5 $\uparrow$5.8 $\uparrow$10.4 $\uparrow$6.6 $\uparrow$5.6 $\uparrow$4.7 $\uparrow$6.2 $\uparrow$16.2 $\uparrow$5.2
InternVL3-8B [53] 34.7 56.9 26.0 31.1 26.1 40.3 27.4 33.1 34.9 39.4 32.2
$\Delta$uparrow7.2 $\uparrow$15.2 $\uparrow$3.8 $\uparrow$6.7 $\uparrow$4.4 $\uparrow$7.3 $\uparrow$6.5 $\uparrow$5.1 $\uparrow$7.6 $\uparrow$10.6 $\uparrow$5.2
Performance biases in VLMs for damage counting. Remote sensing imagery typically encompasses numerous objects exhibiting diverse scales and morphologies, with counting challenges becoming particularly acute when conducting fine-grained damage assessment. Fig. 9 illustrates building damage assessment accuracy as a function of building density within analyzed scenes. For InternVL series models, performance initially declines and then improves as density increases. For peripheral ranges ($<$50 or $\geq$200 buildings), these models demonstrate higher confidence and accuracy. In contrast, GPT-4 models exhibit a clear inverse relationship between building density and accuracy. The fine-tuned InternVL3-8B exhibits substantial improvement in low-density scenes ($<$50 buildings) but notable degradation in other ranges, revealing an overfitting dilemma. Different VLMs have different biases in the damage assessment task. In the future, we can integrate pixel-level semantics provided by the DisasterM3 dataset to alleviate the overfitting risk.
@@
Effects of different prompts. As shown in Fig. 9, we evaluated the robustness of VLMs with five different prompts, where quartiles, ranges of accuracies are plotted. Due to limited LLM capabilities, LLaVA-OV, TeoChat, EarthDial, and Kimi models exhibit higher sensitivity to prompt variations. Besides, IntvernVL3 and Qwen2.5-VL series models show similar patterns wherein larger LLMs display enhanced stability. In comparison to GPT-4o, GPT-4.1 achieves superior performance with notably improved consistency. Following enrichment with the disaster-specific corpus from the DisasterM3 dataset, th fine-tuned Qwen2.5-VL-7B and InternVL3-8B model demonstrate good stability to prompt variations.

### v2

@@ 4.1 Comparative Results
Domain gap for disaster scenarios. Tab. 2 presents performance evaluations on optical-optical settings for QA tasks. As a traditional VLM, LLaVA-1.5 exhibited significant limitations when processing disaster scenes due to the domain gap. By leveraging extensive multi-modal pretraining datasets and implementing the AnyRes architecture, LLaVA-OV demonstrates enhancements in both accuracy and multi-image processing capabilities. As efficient Mixture-of-Experts (MoE) VLMs, Kimi-VL‑A3B‑Think exceeds Kimi-VL‑A3B‑Instruct in mathematical counting tasks (BDC, DRE). However, the non-negligible domain gap limits their application on complex tasks, particularly degrading performance to near-random levels on the ORR task. This motivated our development of the DisasterM3 dataset, which identifies performance gaps through the Bench set while providing complementary training data via the Instruct set.
@@
LLaVA-1.5-7B liu2023visual 12.1 4.2 - - - - 20.0 - - - - - - - -
LLaVA-OV-7B li2024llava 24.5 16.3 53.5 3.7 26.4 24.2 22.7 1.66 1.50 1.53 1.93 2.30 3.01 2.08 1.81
Kimi-VL-A3B-Instruct team2025kimi 25.6 28.9 66.3 4.0 20.4 15.0 18.9 1.69 1.53 1.72 1.81 2.67 3.57 2.40 2.05
Kimi-VL-A3B-Think team2025kimi 26.7 27.0 51.6 7.4 24.4 25.4 24.4 1.61 1.39 1.68 1.75 2.61 3.35 2.34 2.15
InternVL3-8B zhu2025internvl3 31.3 39.6 53.5 4.0 30.3 24.1 36.2 1.96 1.88 1.92 2.09 2.75 3.52 2.53 2.21
InternVL3-14B zhu2025internvl3 35.7 42.5 62.0 4.9 27.4 23.6 54.1 2.08 2.01 2.01 2.22 2.86 3.67 2.62 2.29
InternVL3-78B zhu2025internvl3 39.3 43.5 72.5 5.3 29.4 28.7 56.1 2.79 2.74 2.75 2.89 2.90 3.64 2.64 2.43
Qwen2.5-VL-3B bai2025qwen2 26.2 30.8 56.1 5.7 29.9 21.2 13.8 1.00 0.83 1.05 1.12 2.15 2.98 1.77 1.71
Qwen2.5-VL-7B bai2025qwen2 31.2 28.3 66.6 4.7 34.2 29.3 23.9 1.75 1.69 1.71 1.85 1.95 2.53 1.83 1.49
Qwen2.5-VL-32B bai2025qwen2 35.3 36.7 54.7 11.6 33.2 30.9 44.8 1.55 1.42 1.52 1.72 2.96 3.63 2.71 2.55
Qwen2.5-VL-72B bai2025qwen2 40.5 47.0 74.8 6.8 34.8 28.9 50.8 2.01 1.99 2.00 2.05 2.92 3.79 2.70 2.27
GeoChat-7B kuckreja2024geochat 10.7 6.1 - - - - 15.3 - - - - - - - -
TeoChat-7B irvin2024teochat 23.0 6.9 64.9 2.0 22.5 23.3 18.2 1.77 1.61 1.74 1.96 1.95 2.59 1.77 1.49
EarthDial-4B soni2024earthdial 22.9 10.6 58.1 3.2 30.2 20.8 14.5 1.53 1.22 1.64 1.73 2.42 3.21 2.08 1.98
@@
Qwen2.5-VL-7B bai2025qwen2 40.4 37.7 83.6 21.5 34.3 29.4 36.2 3.90 3.76 3.53 4.41 3.11 3.73 2.88 2.73
$\Delta$ $\uparrow$9.2 $\uparrow$9.4 $\uparrow$17.0 $\uparrow$16.8 $\uparrow$0.1 $\uparrow$0.1 $\uparrow$12.3 $\uparrow$2.15 $\uparrow$2.07 $\uparrow$1.82 $\uparrow$2.56 $\uparrow$1.26 $\uparrow$1.20 $\uparrow$1.83 $\uparrow$1.24
InternVL3-8B zhu2025internvl3 41.7 42.6 79.3 23.9 29.1 24.9 50.6 3.83 3.69 3.49 4.32 3.31 3.92 3.10 2.90
$\Delta$ $\uparrow$10.4 $\uparrow$3.0 $\uparrow$25.8 $\uparrow$19.9 $\downarrow$-1.2 $\uparrow$0.8 $\uparrow$14.4 $\uparrow$1.87 $\uparrow$1.81 $\uparrow$1.57 $\uparrow$2.23 $\uparrow$0.56 $\uparrow$0.40 $\uparrow$0.57 $\uparrow$0.69
Larger VLMs achieve higher performances. By scaling up LLMs, InternVL3 and Qwen2.5-VL series demonstrate consistent trends that larger LLMs achieve superior performances, confirming established scaling laws observed in general-domain applications. The commercial models, i.e., GPT-4o and GPT-4.1, showcase competitive performances across all tasks due to their massive corpus.
@@
Fine-tuned models improve comprehensively. By fine-tuning on DisasterM3 Instruct set, the performances of Qwen2.5-VL and InternVL3 have been significantly improved, narrowing the domain gap. Disaster-specific terminology integration during training significantly enhances report generation quality, resulting in more reasonable and professional reports. However, for building damage counting (BDC) task, the fine-tuned InternVL3 exhibits unexpected performance degradation due to overfitting, and we perform detailed analysis in §4. In the future, object sensitive module zhu2024zero and numerical enhanced optimization wang2024earthvqa could be explored for model development.
@@
Underrepresentation for SAR images. Disasters are often accompanied by extreme weather, with clouds and rain blocking optical sensors. In this case, the active imaging method SAR can penetrate clouds and fog to obtain accurate surface information. Fig. 7 shows the VLMs’ performances evaluated on paired optical-SAR images. Due to the reduced semantics compared to optical imagery and underrepresentation in generic VLM, the performance using post-SAR images yielded substantially diminished performance across all evaluation tasks. In this scenario, as demonstrated by the Qwen2.5-VL series, increasing LLM size fails to yield stable improvements. The fine-tuned models alleviate the multi-sensor feature alignment, significantly improving the performances on all tasks. Since there still exists a huge gap compared to the optical-optical setting, more multi-modal pretraining 10490262 and alignment strategies could be investigated wang2023multi for further improvement.
Table 3: Benchmarking results of referring segmentation VLMs on DisasterM3 Bench set. The accuracies across damage-levels (buildings and roads) are combined for simplification.
@@
PSALM-1.3B zhang2025psalm 9.7 6.3 2.6 10.2 16.3 8.1 8.8 5.1 11.1
HyperSeg-3B wei2024hyperseg 16.6 14.5 7.5 17.0 25.4 8.8 10.3 4.5 13.1
LISA-7B lai2024lisa 27.5 22.1 11.9 25.0 45.6 10.9 12.7 6.0 15.7
GeoPixel-7B shabbir2025geopixel 14.3 14.2 8.5 18.1 16.2 4.0 5.1 1.8 6.2
@@
LISA-7B lai2024lisa 44.8 43.7 27.6 41.2 65.5 28.2 29.6 21.5 34.9
$\Delta$ $\uparrow$17.3 $\uparrow$21.6 $\uparrow$15.7 $\uparrow$16.2 $\uparrow$19.9 $\uparrow$17.3 $\uparrow$16.9 $\uparrow$15.5 $\uparrow$19.2
PSALM-1.3B zhang2025psalm 50.5 44.5 30.5 49.1 71.9 31.8 35.2 24.3 39.3
$\Delta$ $\uparrow$40.8 $\uparrow$38.2 $\uparrow$27.9 $\uparrow$38.9 $\uparrow$55.6 $\uparrow$23.7 $\uparrow$26.4 $\uparrow$19.2 $\uparrow$28.2
@@ 4.2 Detailed Analysis
LLaVA-OV li2024llava 21.2 23.6 17.1 19.2 22.8 25.4 18.1 19.5 23.2 23.5 19.9
Kimi-VL-A3B-Instruct team2025kimi 19.9 22.2 17.5 20.1 16.3 26.3 13.2 19.2 21.9 23.5 18.4
Kimi-VL-A3B-Think team2025kimi 22.0 26.4 19.6 21.0 17.4 26.9 16.9 22.6 21.8 25.0 22.3
InternVL3-8B zhu2025internvl3 27.5 41.7 22.2 24.4 21.7 33.0 20.9 28.0 27.3 28.8 27.0
InternVL3-14B zhu2025internvl3 30.0 48.6 22.7 26.6 27.2 33.7 21.1 27.3 29.9 31.3 31.5
InternVL3-78B zhu2025internvl3 31.8 48.6 26.3 27.9 25.0 37.1 25.6 32.0 31.7 32.9 30.9
Qwen2.5-VL-3B bai2025qwen2 24.5 26.4 19.5 21.9 27.5 32.1 17.9 24.2 24.2 29.6 21.8
Qwen2.5-VL-7B bai2025qwen2 25.6 34.3 21.0 24.9 17.3 33.7 16.7 28.3 27.8 25.6 25.9
Qwen2.5-VL-32B bai2025qwen2 31.0 50.0 26.4 27.1 26.6 35.7 23.9 32.8 29.4 30.8 27.7
Qwen2.5-VL-72B bai2025qwen2 31.8 47.2 25.0 31.1 19.0 39.0 24.0 33.4 34.0 33.9 31.2
GPT-4o hurst2024gpt 30.7 52.8 24.8 25.7 19.6 33.7 25.6 28.0 29.7 35.1 32.0
GPT-4.1 hurst2024gpt 32.4 51.4 26.9 26.7 21.7 35.2 27.7 28.5 33.4 35.8 36.5
@@
Qwen2.5-VL-7B bai2025qwen2 32.9 41.7 26.5 30.7 27.7 40.3 22.3 33.0 34.0 41.8 31.1
$\Delta$ $\uparrow$7.3 $\uparrow$7.4 $\uparrow$5.5 $\uparrow$5.8 $\uparrow$10.4 $\uparrow$6.6 $\uparrow$5.6 $\uparrow$4.7 $\uparrow$6.2 $\uparrow$16.2 $\uparrow$5.2
InternVL3-8B zhu2025internvl3 34.7 56.9 26.0 31.1 26.1 40.3 27.4 33.1 34.9 39.4 32.2
$\Delta$ $\uparrow$7.2 $\uparrow$15.2 $\uparrow$3.8 $\uparrow$6.7 $\uparrow$4.4 $\uparrow$7.3 $\uparrow$6.5 $\uparrow$5.1 $\uparrow$7.6 $\uparrow$10.6 $\uparrow$5.2
Performance biases in VLMs for damage counting. Remote sensing imagery typically encompasses numerous objects exhibiting diverse scales and morphologies, with counting challenges becoming particularly acute when conducting fine-grained damage assessment. Fig. 9 illustrates building damage assessment accuracy as a function of building density within analyzed scenes. For InternVL series models, performance initially declines and then improves as density increases. For peripheral ranges (<<50 or $\geq$200 buildings), these models demonstrate higher confidence and accuracy. In contrast, GPT-4 models exhibit a clear inverse relationship between building density and accuracy. The fine-tuned InternVL3-8B exhibits substantial improvement in low-density scenes (<<50 buildings) but notable degradation in other ranges, revealing an overfitting dilemma. Different VLMs have different biases in the damage assessment task. In the future, we can integrate pixel-level semantics provided by the DisasterM3 dataset to alleviate the overfitting risk.
@@
Effects of different prompts. As shown in Fig. 9, we evaluated the robustness of VLMs with five different prompts, where quartiles, ranges of accuracies are plotted. Due to limited LLM capabilities, LLaVA-OV, TeoChat, EarthDial, and Kimi models exhibit higher sensitivity to prompt variations. Besides, InternVL3 and Qwen2.5-VL series models show similar patterns wherein larger LLMs display enhanced stability. In comparison to GPT-4o, GPT-4.1 achieves superior performance with notably improved consistency. Following enrichment with the disaster-specific corpus from the DisasterM3 dataset, the fine-tuned Qwen2.5-VL-7B and InternVL3-8B model demonstrate good stability to prompt variations.
5 Limitations and Future Directions
While DisasterM3 represents a significant step forward in disaster-oriented vision-language research, we acknowledge several limitations that open avenues for future work. 1) Multi-resolution generalization: Our standardization to 0.8 m resolution ensures controlled experimentation but limits evaluation of model robustness to diverse spatial resolutions encountered in operational settings. Future work should incorporate multi-resolution imagery from platforms like Sentinel-2 (10m) and Landsat (30m), leveraging our existing annotations through geo-registration. 2) Enhanced sensor diversity: Although we include both optical and SAR imagery, our SAR data is limited to single polarization. Integrating multi-polarization data (e.g., Sentinel-1’s VV+VH) would provide richer scattering information about debris orientation and surface characteristics, enabling more comprehensive damage assessment. 3) Cross-sensor performance gap: The significant performance degradation on SAR imagery highlights the need for advanced multi-modal pretraining and cross-modal alignment strategies to better bridge the optical-SAR domain gap. 4) Counting task optimization: To address overfitting in damage object counting, promising directions include object-sensitive encoders (e.g., DINOv2), numerical difference loss, synthetic data generation via diffusion models for high-density scenarios, and knowledge distillation strategies. 5) Living benchmark commitment: We will maintain DisasterM3 as an evolving resource by regularly incorporating new disaster events from the Maxar Open Data Program, ensuring continued relevance and growth in geographic and temporal coverage for the disaster response community.